### v1

@@ 4 Coordination Increases Bidders’ Value
5 Extension to Non-i.i.d. Valuations
We relax the i.i.d. assumption and allow bidder-specific value distributions. In each round $t\in[T]$, bidder $i\in[N]$ draws a value $v_{i,t}\sim F_{i}$ (independently across bidders and rounds), and the outside bid $d^{O}_{t}$ is drawn from $D$. Values and outside bids are bounded by $B>0$. Let $V_{t}^{C}$ and $V_{t}^{I}$ denote the coalition’s total value achieved in round $t$ under coordinated and independent bidding, respectively. Let $U_{i}^{\cdot,A}$ denote the cumulative RoS violation (achieved value minus payment) of bidder $i$ over $T$ rounds under scenario ${\cdot}\in\{C,I\}$ when all bidders follow auto-bidding algorithm $A$.
Total RoS violation.
Our first result shows that, in the non-i.i.d. setting, coordination strictly reduces the total RoS constraint violation, namely, the total utility of the coalition bidders is improved:
Theorem 5.1(Coordination reduces Total RoS Violation, non-iid case)
Assume 1. For any auto-bidding algorithm $A$ that (weakly) overbids, with probability at least $1-\exp\!\big(-\tfrac{\Delta^{2}T}{32B^{2}}\big)$, the total utility $\sum_{i=1}^{N}U_{i}^{C,A}\geq\sum_{i=1}^{N}U_{i}^{I,A}+\frac{\Delta T}{2}$.
See Appendix 0.B for the proof.
Total value.
To compare total value, we require a per-bidder margin condition that ensures each bidder wins with a positive average advantage when they are the highest-value coalition member. For bidder $i$, define $\Delta_{i}:=\mathbb{E}\left[(v_{i}-d^{O})\,\mathbb{I}\{v_{i}=\max_{j\in[N]}v_{j}\}\right]$.
Assumption 2 (Positive per-bidder winning margin)
For every bidder $i\in[N]$, $\Delta_{i}>0$.
Theorem 5.2(Coordination Increases Total Value, non-iid case)
Assume 2, and suppose bidders run Algorithm 3. Then, we have $\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0$.
See Appendix 0.C for the proof.
1(a) i.i.d. 2 4000 -0.011 0.220 0.643 0.666
1(b) i.i.d. 4 4000 -0.077 0.302 0.773 0.800
1(c) i.i.d. 3 4000 -0.048 0.154 0.713 0.749
2(a) non-i.i.d. 2 10000 0.049 0.219 0.715 0.718
2(b) non-i.i.d. 3 20000 -0.014 0.258 0.620 0.633
2(c) non-i.i.d. 5 20000 -0.062 0.619 0.814 0.819
3(a) Real data 4 20000 -0.040 0.156 $\pm$ 0.012 0.614 $\pm$ 0.016 0.929
3(b) Real data 5 20000 -0.065 0.174 $\pm$ 0.012 0.607 $\pm$ 0.012 0.959
Table 1: Performance summary. Confidence intervals below $\pm 0.002$ are omitted.
(a) $N=2,F=U[0,1],D=U[0,0.9]$
(b) $N=4,F=U[0,1],D=U[0,1]$
(c) $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$
Figure 1: Experiments under i.i.d auto-bidders.
6 Experiments
To complement our theoretical analysis, we conduct a series of experiments evaluating the performance of coordinated auto-bidding, on both synthetic data and a public real-world dataset (iPinYou [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou]).
Unless otherwise stated, all metrics are averaged over 100 independent simulations; shaded regions in figures depict 95% confidence intervals computed as $\bar{x}\pm 1.96\cdot s$, where $\bar{x}$ and $s$ are the sample mean and standard deviation over the 100 simulations.
6.1 Experiment Setup
For all experiments, we run Algorithm 3 with learning rate $\alpha=1/\sqrt{T}$, where $T$ is the total number of auction rounds. The horizon $T$ may vary across experiments because different settings exhibit different convergence rates. The values of $T$ and $N$ (number of coalition bidders) used in each setting are summarized in Table 1.
6.1.1 Synthetic Data
Symmetric setting.
Each coalition bidder $i$ draws values i.i.d. from uniform distribution $U[0,1]$. The outside bid is sampled from $D=U[0,0.9]$ (Figure 1(a)), $D=U[0,1]$ (Figure 1(b)), or $D=\mathrm{Beta}(3,2)$ (Figure 1(c)), representing different market conditions faced by the coalition bidders.
Asymmetric setting.
To stress-test the mechanism under heterogeneity, we sample each bidder’s values from distinct families (Uniform, Beta, or truncated Gaussians with bidder-specific parameters), as summarized in Figure 2. The outside-bid distribution $D$ varies with the scenario: $U[0.2,0.8]$ in Figure 2(a), $\mathrm{Beta}(3,5)$ in Figure 2(b), and $\mathrm{Beta}(2,8)$ in Figure 2(c). Asymmetries arise from both inside the coalition and the outside market conditions.
Full distributional specifications for the synthetic experiments are provided in the figure captions (Figures 1 and 2). All distributions conform to Assumption 1.
6.1.2 Real-World Data
We also experimented with the public iPinYou dataset [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou], released for a real-time bidding competition organized by iPinYou. We use the Season 2 test split, which contains 2,521,6302{,}521{,}630 auction records from $5$ advertisers. Each record provides the bidding (winning) price and the corresponding advertiser ID.
The real-world data impose three limitations: (i) only the winning price is observed, (ii) bidder identities are unavailable, and (iii) there is no label indicating whether any bidders coordinate. To obtain bidder valuations consistent with the data while abstracting away unobserved identities and strategies, we adopt an empirical i.i.d. model: we aggregate all observed winning prices into a common pool and, in each round, draw each coalition bidder’s value independently and uniformly at random with replacement from this pool. This is equivalent to sampling i.i.d. values from the empirical price distribution induced by the dataset.
Following prior work [chen_coordinated_2023], we normalize all prices to $[0,1]$ for numerical stability. The outside bidder’s bid is drawn from the same empirical distribution and then scaled by a random factor from $U[1,2]$ to maintain competitive pressure. We report results for coalitions with $N\in\{4,5\}$ bidders under this setup.
6.2 Results
Metrics.
Across all experiments, we evaluate three metrics for both Independent (Algorithm 1) and Coordination (Algorithm 2): (a) utility (achieved value minus payment), (b) achieved value, and (c) multiplier. Higher utility indicates lower RoS violation; in particular, utility $\geq 0$ implies that the RoS constraint is satisfied. Figures report per-bidder trajectories, whereas Table 1 aggregates to coalition-level totals (sum across bidders). All statistics are averaged over 100 independent runs.
Plot notation.
In the figures, labels of the form bidder $i$ -- I denote bidder $i$ under the Independent scenario, whereas bidder $i$ -- C denotes the Coordination scenario. Shaded regions show 95% confidence intervals over 100 simulations.
Results on synthetic data
In the symmetric settings (Figure 1), coordination sharply increases both each bidder’s utility (hence reducing RoS violation) and the coalition’s total achieved value relative to independent bidding. Although our theory guarantees value improvement at the coalition (total) level, the experiments further show that every individual bidder benefits from coordination by obtaining higher value. Moreover, all multipliers converge to 0, consistent with our theory, and per-bidder statistics (violation, value, and multiplier trajectories) closely track one another due to the i.i.d. value draws.
In the asymmetric settings (Figure 2), the aggregate gains largely persist: coordination improves total value on average and typically improves violations. However, heterogeneity induces dispersion in learning dynamics – bidders with different value distributions exhibit different convergence rates of the multiplier, and individual value improvements are not guaranteed for every bidder even though the coalition-level total value increases. In cases where each bidder has a non-negligible chance to be the winning bidder, we also observe that the multipliers tend to converge toward 0 empirically.
Results on real-world data.
Consistent with the synthetic data experiments, real-world data results show improvements in both individual RoS violation and achieved value, accompanied by multipliers converging steadily toward 0.
7 Discussion
In this paper, we studied coordinated auto-bidding mechanisms for value maximization under RoS constraints, to understand how cooperation among learning agents enhances efficiency in repeated auctions. We proposed a simple yet effective coordination protocol and theoretically proved that coordination improves the coalition’s total value, as well as RoS constraint compliance, under mild assumptions. Our empirical results further validate these findings, showing higher achieved value and consistent convergence patterns. The demonstrated benefit of coordination over independent bidding might provide useful guidance for industrial practice.
  • Auto-bidding outside bidders: While we assumed i.i.d. outside bids (a reasonable assumption when outside bidders are uniformly sampled from a population every round), a natural extension is to incorporate auto-bidding outside bidders, allowing both coalition members and external participants to adjust their bids dynamically.
  • Alternative coordination mechanisms: Another direction is to explore alternative coordination mechanisms, such as highest-bid coordination, to examine how they would affect the bidders’ overall welfare. Highest-bid coordination is analogous to highest-value coordination in the symmetric setting, but not in the asymmetric setting.
Appendix 0.A Proof of Lemma 3
$G(\lambda;v)$ $=\mathbb{E}_{d^{O}\sim D}\big[(v-d^{O})\cdot\mathbb{I}[(1+\tfrac{1}{\lambda})v>d^{O}]\big]$
$=vx\big((1+\tfrac{1}{\lambda})v\big)-p\big((1+\tfrac{1}{\lambda})v\big),$
where $x(b)$ is the winning probability when the bidder bids $b$ and $p(b)$ is the expected payment when the bidder bids $b$. Since the second-price auction is a truthful auction, by Myerson’s payment identity for truthful auctions, we have
$p(b)=xb-\int_{0}^{b}x(z)\mathrm{d}z.$ (7)
So,
$G(\lambda;v)$
$=vx\big((1+\tfrac{1}{\lambda})v\big)-(1+\tfrac{1}{\lambda})vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz$
Take derivative with respect to $\lambda$:
$\tfrac{\partial G(\lambda;v)}{\partial\lambda}$
$=\tfrac{1}{\lambda^{2}}vx\big((1+\tfrac{1}{\lambda})v\big)-\tfrac{1}{\lambda}vx^{\prime}\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})+x\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})$
$=\tfrac{v^{2}}{\lambda^{3}}x^{\prime}\big((1+\tfrac{1}{\lambda})v\big)$
$\geq 0$
where the last inequality is because $x^{\prime}(b)\geq 0$. Integrating over $v\sim F_{(N)}$, we obtain $G^{\prime}(\lambda)\geq 0$, hence $G(\lambda)$ is increasing.
Appendix 0.B Proof of Theorem 5.1
@@ Appendix 0.A Proof of Lemma 3
Let $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$ denote the index of the highest-value bidder in round $t$. Conditioned on the history $H_{t-1}$, we analyze the expected increment $\Delta_{i,t}=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$ for bidder $i$:
$\mathbb{E}_{\bm{v}_{t}\sim F,\,d_{t}^{O}\sim D}\!\left[\Delta_{i,t}\mid H_{t-1}\right]$
$=\mathbb{E}_{\bm{v}_{t},\,d_{t}^{O}}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\Delta_{i,t}\mid H_{t-1}\right]$
$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
$\geq\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\,\big(\max_{j\neq i}v_{j,t}-d_{t}^{O}\big)_{+}\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}]\cdot(v_{i,t}-d_{t}^{O})\right].$
@@
where the second inequality follows from Assumption 1. Finally, applying Azuma’s inequality (in a similar way as in the proof of Theorem 3.1) completes the proof.
Appendix 0.C Proof of Theorem 5.2
@@
Let $G_{i}^{C}(\lambda_{i})$ denote bidder $i$’s expected utility when $v_{i}\sim F_{i}$ and the coalition uses multiplier $\lambda_{i}>0$:
@@
By reusing the proof of Lemma 3 (Appendix 0.A) with $D$ replaced by the distribution of $\max\{d^{O},\max_{j\neq i}v_{j}\}$, we claim that $G_{i}^{C}(\lambda_{i})$ is increasing in $\lambda_{i}$. Hence,
$G_{i}^{C}(\lambda_{i})\geq\lim_{\mu\to 0^{+}}G_{i}^{C}(\mu)=\mathbb{E}\!\left[(v_{i}-d^{O})\,\mathbb{I}\!\left\{v_{i}=\max_{j\in[N]}\{v_{j}\}\right\}\right]=:\Delta_{i},$
@@
We now present the parallel version of Lemma 5.
Lemma 7
@@
$\lambda_{i,t+1}^{C}\;\leq\;\lambda_{i,1}\,\exp\!\Big(-\frac{\Delta_{i}\,t}{2\sqrt{T}}\Big).$
Proof
@@
Applying Azuma’s inequality (exactly as in the proof of Lemma 5) completes the argument. ∎
@@
$V_{(N)}(\lambda)$ $\;\coloneqq\;\mathbb{E}_{\,v_{(N)}\sim F^{\max},\,d^{O}\sim D}\Big[\,v_{(N)}\cdot\mathbb{I}\big\{(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big\}\Big],$
$V_{(N)}(0^{+})$ $\;\coloneqq\;\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda).$
@@
In the non-i.i.d. case, Lemma 6 still holds:
Lemma 8(copy of Lemma 6)
$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$.
Proof(Proof of Theorem 5.2)
@@
$\Delta_{\min}\coloneqq\min_{i\in[N]}\Delta_{i}>0,\qquad c_{\min}\coloneqq\frac{\Delta_{\min}^{2}}{32B^{2}}$
Define
$m_{T}\;\coloneqq\;\Big\lceil\frac{2\sqrt{T}}{\Delta_{\min}}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\Big\rceil.$
By Lemma 7, for each $i$ and every $t\geq m_{T}$, $\lambda_{i,t+1}^{C}\leq\delta$ holds with probability at least $1-\exp(-c_{\min}t)$. A union bound over $i=1,\dots,N$ and $t=m_{T},\dots,T$ yields
@@
$\mathbb{E}\Big[\Big|V_{(N)}(0^{+})-\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\Big]\;\leq\;\frac{m_{T}}{T}B+\varepsilon+B\,p_{T}\;\xrightarrow[T\to\infty]{}\;0.$
@@
$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]=V_{(N)}(0^{+}).$
Finally, combining with Lemma 8 yields the desired inequality against the independent-bidding benchmark, completing the proof. ∎
Appendix 0.D Additional Experiment Results
This appendix presents additional figures in the experiments: Figure 2 for the asymmetric setting on synthetic data; Figure 3 for the real-world dataset.
(a) $F_{1}=\mathrm{Beta}(2,5),F_{2}=\mathrm{Beta}(5,2),D=U[0.2,0.8],N=2$
(b) $F_{1}=\mathrm{Beta}(2,6),F_{2}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.6,0.15),F_{3}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.4,0.2),D=\mathrm{Beta}(3,5),N=3$
(c) $F_{1}=U[0.2,0.8],F_{2}=\mathrm{Beta}(4,3),F_{3}=\mathrm{Beta}(6,2),F_{4}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.5,0.1),F_{5}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.7,0.12),D=\mathrm{Beta}(2,8),N=5$

### v2

@@ B.2 Boundary case $\lambda_{\star}=0$
Appendix C Omitted Proofs in Section 5
C.1 Proof of Theorem 5.1
@@ Appendix C Omitted Proofs in Section 5
Let $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$ denote the index of the highest-value bidder in round $t$. Conditioned on the history $H_{t-1}$, we analyze the expected difference $\Delta_{i,t}=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$ for bidder $i$:
$\mathbb{E}_{\bm{v}_{t}\sim F,\,d_{t}^{O}\sim D}\big[\Delta_{i,t}\mid H_{t-1}\big]$
$=\mathbb{E}_{\bm{v}_{t},\,d_{t}^{O}}\big[\mathbb{I}[i^{*}_{t}=i]\cdot\Delta_{i,t}\mid H_{t-1}\big]$
$=\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\Big]+\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\Big]$
$=\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\Big]+\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\Big]$
$\geq\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\,\big(\max_{j\neq i}v_{j,t}-d_{t}^{O}\big)_{+}\Big]+\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}]\cdot(v_{i,t}-d_{t}^{O})\Big].$
@@
where the second inequality follows from Assumption 1. Summing over $t\in[T]$ gives $\mathbb{E}[\sum_{i=1}^{N}U_{i,T}^{C,A}-\sum_{i=1}^{N}U_{i,T}^{I,A}]\geq\Delta T$.
C.2 Proof of Theorem 5.2
@@
Let $G_{i}^{C}(\lambda_{i})$ denote bidder $i$’s expected utility when $v_{i}\sim F_{i}$ and bidder $i$ uses multiplier $\lambda_{i}>0$ in the coordinated scenario:
@@
By reusing the proof of Lemma B.1 with $D$ replaced by the distribution of $\max\{d^{O},\max_{j\neq i}v_{j}\}$, we claim that $G_{i}^{C}(\lambda_{i})$ is increasing in $\lambda_{i}$. Hence,
$G_{i}^{C}(\lambda_{i})~\geq~\lim_{\mu\to 0^{+}}G_{i}^{C}(\mu)~=~E\Big[(v_{i}-d^{O})\,\mathbb{I}\big\{v_{i}=\max_{j\in[N]}\{v_{j}\}\big\}\Big]~=:~\Delta_{i},$
@@
We now present the parallel version of Lemma B.7.
Lemma C.1.
@@
$h^{\prime}(\lambda_{i,t+1}^{C})\leq h^{\prime}(\lambda_{i,1})-\frac{\Delta_{i}t}{2\sqrt{T}}.$
Proof.
@@
Applying Azuma’s inequality (exactly as in the proof of Lemma B.7) completes the argument. ∎
@@
$V_{(N)}(\lambda)$ $~\coloneqq~\mathbb{E}_{\,v_{(N)}\sim F^{\max},\,d^{O}\sim D}\Big[\,v_{(N)}\cdot\mathbb{I}\big\{(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big\}\Big],$
$V_{(N)}(0^{+})$ $~\coloneqq~\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)~=~\mathbb{E}_{\,v_{(N)}\sim F^{\max}}\big[\,v_{(N)}\big].$
@@
In the non-i.i.d. case, Lemma B.8 still holds:
Lemma C.2 (copy of Lemma B.8).
For any coordination mechanism $\mathcal{G}$ and any round $t$,
$\mathbb{E}\!\left[V_{t}^{\mathcal{G}}\right]~\leq~V_{(N)}(0^{+})~=~\mathbb{E}_{v_{(N)}\sim F^{\max}}\big[v_{(N)}\big].$
Consequently, for every $T\geq 1$,
$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{\mathcal{G}}\Big]~\leq~V_{(N)}(0^{+}).$
Proof of Theorem 5.2.
@@
$\Delta_{\min}\coloneqq\min_{i\in[N]}\Delta_{i}>0\qquad c_{\min}\coloneqq\frac{\Delta_{\min}^{2}}{32B^{2}}>0.$
Define $b(\delta):=(h^{\prime})^{-1}(\delta)$ and
$m_{T}\;\coloneqq\;\Big\lceil\frac{2\sqrt{T}}{\Delta_{\min}}(h^{\prime}(\lambda_{i,1})-b(\delta))\Big\rceil.$
By Lemma C.1, for each $i$ and every $t\geq m_{T}$, $\lambda_{i,t+1}^{C}\leq\delta$ holds with probability at least $1-\exp(-c_{\min}t)$. A union bound over $i=1,\dots,N$ and $t=m_{T},\dots,T$ yields
@@
$\mathbb{E}\Big[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\Big]\;\leq\;\frac{m_{T}}{T}B+\varepsilon+B\,p_{T}\;\xrightarrow[T\to\infty]{}\;0.$
@@
$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]~=~V_{(N)}(0^{+}).$
Finally, combining with Lemma C.2 yields the desired inequality against the independent-bidding total value, completing the proof. ∎
Appendix D Additional Experiment Results
This appendix presents additional figures in the experiments: Figure 3 for the symmetric setting on synthetic data; Figure 4 for the asymmetric setting on synthetic data; Figure 5 for the real-world dataset.
(a) $N=4,F=U[0,1],D=U[0,1]$
(b) $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$
Figure 3: Experiments under i.i.d auto-bidders.
(a) $N=3,F_{1}=\mathrm{Beta}(2,6),F_{2}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.6,0.15),F_{3}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.4,0.2),D=\mathrm{Beta}(3,5)$
(b) $N=5,F_{1}=U[0.2,0.8],F_{2}=\mathrm{Beta}(4,3),F_{3}=\mathrm{Beta}(6,2),F_{4}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.5,0.1),F_{5}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.7,0.12),D=\mathrm{Beta}(2,8)$